Tutorial Exercise Evaluate the expression. C(9,7). C(6,4) C(12,8)
Answer #1

Expression, C09,7).C(6,4) Expression, C(12,8) Caner) = n! (n-1)!r! =36 ((9,7)= 9! (9-7)! 7! C(6,4) = 6! _ =15 (6-4)! 4! = 15